OpenStudy (anonymous):

Please Check these? I don't want to go through the trouble of posting EVERYTHING I wrote here again. I already answered them, I just need someone to say "Yes or No" to if I'm right or not. That is all. Please? Thank You !!

OpenStudy (phi):

the x-3 moves the graph to the right, not the left

OpenStudy (anonymous):

It is up 4. but i think it goes right 3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

because its negative

OpenStudy (anonymous):

lol I know I had to come up with one not so common . :D

OpenStudy (anonymous):

Did they give you a picture

OpenStudy (anonymous):

k will move to the left

OpenStudy (phi):

yes. If you forget that the x is counter-intuitive (backwards from what you might guess) you can look at 2^x with x=0 you get 2^0=1 with 2^(x-3) x-3 is zero when x=3 so 2^(x-3) at x=3 matches 2^x at x=0 it moved right (x moved from 0 to 3)

OpenStudy (anonymous):

no take that back it should be +k

OpenStudy (anonymous):

it can move up or down depending on what the equation would be.

OpenStudy (anonymous):

yes its true it can go both ways

OpenStudy (phi):

Without knowing where the k is in the equation, I can only guess If the equation is y= f(x+h) + k then the k moves up and down.

OpenStudy (phi):

your answer looks ok

OpenStudy (phi):

though if you type it, is should look like this y = -2^(x+1) - 1 parens are important and the ^ means raise to the power

OpenStudy (phi):

yes

OpenStudy (anonymous):

No Problem :D